OpenStudy (anonymous):

Write the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3).

OpenStudy (phi):

parallel lines have equal slopes. What is the slope of y=7x+2 ?

OpenStudy (anonymous):

The slope is 7.

OpenStudy (anonymous):

Sorry for the late reply, hope u can still help.

OpenStudy (phi):

yes. m= 7 so you have (so far) y = 7 x +b now you need to find b. the easy way is plug in (x,y) pair that is on the line. Luckily they tell us (1,-3) is on the line.

OpenStudy (anonymous):

Yes, how do you find b? Do you plug -3 into y?

OpenStudy (phi):

you use both x=1 and y= -3 in y=7x+b then you solve for b.

OpenStudy (anonymous):

I got b = -10.

OpenStudy (phi):

-3= 1*7+b or -3= 7+b to get b by itself, subtract 7 from both sides (or add -7 to both sides) b= -10

OpenStudy (phi):

so the equation is y= 7x-10

OpenStudy (anonymous):

ohh okay! thank you so much @phi

OpenStudy (phi):

Now try to remember all the details, and you can solve all of these type problems.

OpenStudy (anonymous):

Okay.